# Gentle Object Retraction in Dense Clutter Using Multimodal Force Sensing and Imitation Learning

### v1

@@ Abstract
Dense collections of movable objects are common in everyday spaces—from cabinets in a home to shelves in a warehouse. Safely retracting objects from such collections is difficult for robots, yet people do it easily, using non-prehensile tactile sensing on the sides and backs of their hands and arms. We investigate the role of such sensing for training robots to gently reach into constrained clutter and extract objects. The available sensing modalities are (1) “eye-in-hand” vision, (2) proprioception, (3) non-prehensile triaxial tactile sensing, (4) contact wrenches estimated from joint torques, and (5) a measure of successful object acquisition obtained by monitoring the vacuum line of a suction cup. We use imitation learning to train policies from a set of demonstrations on randomly generated scenes, then conduct an ablation study of wrench and tactile information. We evaluate each policy’s performance across 40 unseen environment configurations. Policies employing any force sensing show fewer excessive force failures, an increased overall success rate, and faster completion times. The best performance is achieved using both tactile and wrench information, producing an 80% improvement above the baseline without force information.
@@ I Introduction
Cluttered and constrained environments, like densely packed shelves and cabinets, appear commonly in household and commercial settings. People interact with them daily to place, organize, and extract objects without conscious thought using a combination of tactile sensing and limited vision [1, 2]. In such scenarios, there are often no collision-free paths to reach and extract desired objects. Hence, if robots are to operate in these scenarios, they need to embrace contacts, rather than avoid them, as in traditional motion planning [3]. At the same time, robots need to be mindful of interaction forces to avoid damage.
@@
We provide an investigation of the relative importance of force-torque and tactile sensing for training imitation learning policies to perform gentle, contact-rich manipulation in constrained clutter. We demonstrate that wrench and tactile information provide non-overlapping performance improvements to robots while performing these tasks. To the authors’ knowledge, this work also enables object extraction from more densely cluttered lateral access scenes than prior work (up to 5252% area occupied).
@@ II Related Work
Investigating the importance of force sensing in multimodal systems for contact-rich tasks is an active area of research. Efforts include representation learning of wrench or tactile information [4, 5], using kinesthetic demonstrations[6, 7, 8] or hand-held, robot free demonstrations[9, 10] for tasks that are hard to teleoperate, and learning to perform bimanual manipulation with visuotactile sensors[11, 12]. These approaches provide investigations of the utility of force sensing for performing dexterous, contact-rich manipulation tasks, but do not explore specifically the problem of highly cluttered environments with contact on non-prehensile surfaces.
Other works perform decluttering or packing tasks on tabletop clutter[13, 14, 15, 16, 17], with some approaches using force sensing as part of their multimodal systems[18, 19]. However, in these works the robot can view the manipulation scenario from the top (similar to the overhead view in Fig. 1c), which does not reflect the visual occlusion encountered in lateral access scenarios. Furthermore, an open tabletop environment avoids the sporadic object jamming that frequently occurs in lateral access scenes where the objects are in an enclosed space.
Several recent efforts enable object retraction from cluttered lateral access scenes by performing mechanical search using a slender end-effector with a suction cup at the end [20, 21, 22, 23], yet these approaches are limited to relatively low clutter density, in part due to the lack of force sensing. Specifically, they do not use non-prehensile tactile sensors, despite the majority of manipulation maneuvers being non-prehensile.
Non-prehensile tactile sensors have been shown to enable object motion category classification in constrained clutter[24], but it is non-trivial to incorporate these motion categories into a motion planning system. Prior work has also enabled blind reaching into cluttered environments by leveraging non-prehensile tactile sensing to reach toward target locations amidst fixed obstacles while minimizing contact forces[25, 26, 27, 28]. This approach, however, limits scene variability and avoids the issue of sporadic changes to object impedance when objects jam in constrained yet movable clutter. One recent work uses non-prehensile tactile information to reach toward target locations in constrained, lateral access scenarios by employing motion primitives to react to contacts[29]. All of these approaches assume that target locations are known a priori and do not support target object extraction. Blind object recognition and retraction have been demonstrated using non-prehensile tactile sensing[30], but this is performed in unconstrained tabletop clutter and does not investigate the interplay of vision and force sensing.
@@ III-A Hardware
In addition to camera and tactile information, we provide proprioception readings using the robot’s joint encoders to form a quaternion representation of the robot’s tool coordinate pose, which is located at the suction cup. We also use built-in Flexiv commands that leverage joint-torque sensing to determine the dynamics-compensated wrench, represented in the tool coordinate frame. Considering the force components of this wrench (moments are small at the suction cup), the minimum signal change that can reliably be detected above noise due to motion of the arm is $\approx 3.3\,N$.
@@ III-B Training Setup
Fig. 3 shows the training and inference pipeline used in our system. The $128\times 128$ pixel “eye-in-hand” camera images are fed through a designated ResNet-18 vision encoder, while the $20\times 5$ tactile images are fed through a separate ResNet-18 encoder. ResNet was found to be the most effective pre-trained visual encoder for use on tactile information in prior work [5]. After normalization, pose, pressure, and wrench are concatenated and fed as separate, low-dimensional inputs to the diffusion policy framework. Diffusion policy then outputs actions that can be executed by the robot controller.
@@ IV-B Task Setup
The task involves starting at a home position in front of the cabinet shelf with a goal of gently navigating through the cluttered environment to acquire a target object and bring it out of the shelf. Successful completion of the task requires locating the correct target object (the only red object in the scene), retracting it to the home position, and dropping it into the region outlined in black (visible beneath the suction cup in the foreground of Fig. 4c,d). We require that the task is completed within 120 seconds.
@@
where $F_{net}$ is the L-2 norm of the first 3 elements of the measured wrench and $\delta t_{react}$ is the time given to both demonstrator and policy to react to the contact. We also require that motions cannot exceed a peak force impulse, $I_{peak}$
@@
where $F_{peak}$ is the maximum L-2 norm of the triaxial tactile forces across all taxels.
We implement two distinct impulse thresholds because the peak force is the metric most likely to damage an object, but the wrench measurements are unaware of individual contact magnitudes when there are multiple simultaneous contacts. The peak tactile force, however, does not have full coverage of the robot arm and will be unaware of some contacts, thus a net force threshold is also needed.
@@
Through this process, we establish the excessive impulse task constraints of $I_{net}<20.8\,Ns$ and $I_{peak}<4.8\,Ns$.
@@ IV-C Demonstration Setup
Using a 3Dconnexion SpaceMouse® Compact to teleoperate the robot, we collect demonstrations on randomized cluttered scenes generated through the process described in Section IV-A. As seen in Fig. 5, the demonstrator has access to the “eye-in-hand” camera view, the tactile images as described in Fig. 2, and pressure sensing, as well as an additional “excessive force” image which visually encodes whether the robot is approaching $F_{net}$ and $F_{peak}$. The left and right sides of this excessive force image (shown in Fig. 5b) represent peak tactile force from the left and right sides of the end-effector while the center portion corresponds to the net force magnitude computed using the joint torques. In these portions, we display grayscale values ranging from black to white as the corresponding force values range from 0% to 8080% of $F_{net}$ and $F_{peak}$. Beyond 8080%, we display a red warning in the corresponding portion to enable the demonstrator to respond. This visualization system was empirically tuned to provide the demonstrator with enough warning to avoid excessive force failures while avoiding false positives as necessary non-prehensile maneuvers are performed.
@@ V-A Force Ablation Experiment
With these demonstrations and the architecture described in Section III-B, we trained four policies with varied access to wrench and tactile information to investigate their importance for this task: 1) the baseline policy masks out (replaces with zero values) wrench and tactile information, 2) the wrench-informed policy masks out tactile information, 3) the tactile-informed policy masks out wrench information, and 4) the force-informed policy has both wrench and tactile information included. Apart from this masking, all policies are trained using the same architecture and hyper-parameters, with each policy trained for $200$ epochs.
@@ V-B Results
The results of $40$ evaluations on each policy are summarized in Fig. 6. The first item of note is that each policy that has access to wrench or tactile information in addition to vision significantly outperforms the baseline policy. This can be seen in Fig. 6a, which displays the number of successes, time-out failures, and excessive force failures for each policy. The best performance is demonstrated by the force-informed policy, where we see an 8080% relative improvement above the baseline. The improved performance for all policies with access to either force modality appears to be largely due to a substantial reduction to excessive force failures.
@@
Figure 6: Performance results for 40 evaluations of the force ablation experiment. a) Proportions of successes, time out failures, and excessive force failures for each of the policies. * indicates $p<0.05$ in a two-sided pairwise z-test against the baseline. indicates $p<0.05$ in a two-sided pairwise z-test against the force-informed policy. b) Success times for each policy with error bars showing standard error. Each policy with any force modality improved success rate and accomplished the task faster.
@@
The force-informed policy appears to derive its improved performance largely from the ability to avoid time-out failures as seen in Fig. 6a. In contrast, the significantly higher proportion of timeout failures for the remaining policies often occurred due to pulling away from objects in the absence of large contact forces, failing to focus on the target after it has been isolated, and even retracting without the target object. The baseline policy exhibits the highest timeout failure rate.
Finally, as seen in Fig. 6b, all policies informed by wrench or tactile information accomplished the task faster, with the wrench-informed policy being fastest.
@@
Figure 7: Heatmaps of resulting motion after observing net and peak forces for both the baseline and force-informed policies. a) Motion performed by the baseline policy as a function of observed net forces. b) Resulting motion versus observed peak forces for the baseline policy. c) Resulting motion versus observed net forces for the force-informed policy. d) Resulting motion versus observed peak forces for the force-informed policy. The black dashed lines represent the excessive force values used for our task. The baseline policy displays a higher proportion of large force, low motion datapoints than the force-informed policy, indicating a failure to react to contacts.
@@
To gain a better understanding of how the policies have learned to react to contact forces, we extract the observed forces and resulting displacements of the end-effector for all timesteps of all evaluations of the baseline and force-informed policies. We compute the resulting displacement over a subsequent duration of 0.8s0.8\,s to allow for 1 full inference cycle to enable proper reactions.
@@
The observed motion versus net forces and peak tactile forces can be seen in Fig. 7, where cells toward the bottom right of each plot correspond to a large observed force but little resulting motion—indicating a failure to respond. For reference, the values of $F_{net}$ and $F_{peak}$ are plotted as vertical lines. Ideally, there should be no cells that correspond to little or no motion when forces exceed or approach these values, since remaining in place would maintain the excessive force and cause an impulse failure. As seen in Fig. 7a,b, the baseline policy contains a substantial number of high force, low motion datapoints, but the force-informed policy in Fig. 7c,d avoids low motion as the observed forces increase.
@@
We also note the number of failed reactions by tallying the total instances where less than 0.5cm0.5\,cm of displacement is seen after observing either a net force that exceeds $F_{net}$ or a peak tactile force that exceeds $F_{peak}$. The baseline policy fails to react to large contact forces 412 times in contrast to the force-informed policy seeing only 22 failed reactions.
@@ VI Discussion
When attempting to accomplish object retraction in constrained clutter using imitation learning, we have demonstrated that the addition of force sensing modalities intuitively improves a policy’s ability to be gentle. Beyond this, we note that each policy in our experiment which lacked any force modality that the demonstrator used to react saw a significant increase in timeout failures. These timeout failures appear to be the result of an inability to correctly identify the cause-effect relationship between observations and expert actions.
Previous work establishes a framework to interpret causal relationships in imitation learning by supposing that there are disentangled factors of the observation data, some of which are true causes for expert actions and some of which are nuisance factors that coincidentally correlate with expert actions[34]. This prior work demonstrates that the addition of information—specifically looking at information about the history of actions—can harm policy performance since there are strong correlations between prior actions and expert decisions, but the true causes are due to current observations (i.e. the additional information in the history contains a significant amount of nuisance factors). In this way, removal of nuisance factors from the observation data can reduce causal confusion in imitation learning.
We suspect that in our experiments, the baseline policy has overfit to nuisance factors in the absence of force information, which was the true cause of demonstrator reactions. For example, consider the instances where a quick motion to the right was performed after the demonstrator observed a large peak tactile force on the left. Further suppose that in many of these instances, a specific face of the yellow objects used in our scenes coincidentally appeared in the top right of the camera images leading up to these reactions—thus, we have a nuisance factor in our observation data. During rollouts, when this visual observation reappears and no longer correlates to a large force on the left, a policy which lacks the true, force-informed causes of such reactions is more susceptible to the nuisance factor and performs an inappropriate reaction to the right. The accumulation of such reactions may be the cause of the increased timeout failures for all policies which lacked any force modality used by the demonstrator. This indicates that causal confusion can be remedied not only by removing observation data that is ridden with nuisance factors but also by adding observation data that contains true causes.

### v2

@@ Abstract
Dense collections of movable objects are common in everyday spaces—from cabinets in a home to shelves in a warehouse. Safely retracting objects from such collections is difficult for robots, yet people do it frequently, leveraging learned experience in tandem with vision and non-prehensile tactile sensing on the sides and backs of their hands and arms. We investigate the role of contact force sensing for training robots to gently reach into constrained clutter and extract objects. The available sensing modalities are (1) “eye-in-hand” vision, (2) proprioception, (3) non-prehensile triaxial tactile sensing, (4) contact wrenches estimated from joint torques, and (5) a measure of object acquisition obtained by monitoring the vacuum line of a suction cup. We use imitation learning to train policies from a set of demonstrations on randomly generated scenes, then conduct an ablation study of wrench and tactile information. We evaluate each policy’s performance across 40 unseen environment configurations. Policies employing any force sensing show fewer excessive force failures, an increased overall success rate, and faster completion times. The best performance is achieved using both tactile and wrench information, producing an 80% improvement above the baseline without force information.
@@ I Introduction
Cluttered and constrained environments, like densely packed shelves and cabinets, appear commonly in household and commercial settings. People interact with them daily to place, organize, and extract objects using a combination of tactile sensing and limited vision [1, 2]. Visual information aids in target object identification and provides an initial sense of the obstacle configuration while tactile information can inform when objects are immovable and improve object localization when visual occlusions are present during reaching. In such cluttered scenarios, there are often no collision-free paths to reach and extract desired objects. Hence, if robots are to operate in these scenarios, they need to embrace contacts, rather than avoid them, as in traditional motion planning [3]. At the same time, robots need to be mindful of interaction forces to avoid damage.
@@
We provide an investigation of the relative importance of force-torque and tactile sensing for training imitation learning policies to perform gentle, contact-rich manipulation in constrained clutter. We show that wrench and tactile information individually provide performance improvements to robots while performing these tasks, but we see additional improvements when using both modalities. This work demonstrates object extraction from densely cluttered lateral access scenes (up to 5555% area occupied). While clutter density is an important factor determining task difficulty, additional factors such as the shape variability, deformability, and optical or spectral variability are not explicitly investigated.
@@ II Related Work
Utilizing force sensing in multimodal systems for contact-rich tasks is an active area of research. Efforts include representation learning of wrench or tactile information [4, 5], using kinesthetic demonstrations[6, 7, 8] or hand-held, robot free demonstrations[9, 10] for tasks that are hard to teleoperate, and learning to perform bimanual manipulation with visuotactile sensors[11, 12]. These approaches provide investigations of the utility of force sensing for performing dexterous, contact-rich manipulation tasks, but do not explore specifically the problem of highly cluttered environments with contact on non-prehensile surfaces.
Other works perform decluttering, object retraction, or packing tasks on tabletop clutter[13, 14, 15, 16, 17, 18], with some approaches using force sensing as part of their multimodal systems[19, 20]. However, in these works the robot can view the manipulation scenario from the top (similar to the overhead view in Fig. 1c), which does not reflect the visual occlusion encountered in lateral access scenarios. Furthermore, an open tabletop environment avoids the sporadic object jamming that frequently occurs in lateral access scenes where the objects are in an enclosed space.
Several recent efforts enable object retraction from cluttered lateral access scenes by reasoning about occlusions and performing mechanical search to isolate target objects [21, 22, 23, 24, 25], yet these approaches are limited to relatively low clutter density, in part due to the lack of force sensing. Specifically, despite the majority of manipulation maneuvers being non-prehensile, they do not use non-prehensile tactile sensors, which we define as distributed contact force sensors that detect non-prehensile manipulation interactions.
Non-prehensile tactile sensors have been shown to enable object motion category classification in constrained clutter[26], but it is non-trivial to incorporate these motion categories into a motion planning system. Prior work has also enabled blind reaching into cluttered environments by leveraging non-prehensile tactile sensing to reach toward target locations amidst fixed obstacles while minimizing contact forces[27, 28, 29, 30]. This approach, however, limits scene variability and avoids the issue of sporadic changes to object impedance when objects jam in constrained yet movable clutter. One recent work uses non-prehensile tactile information to reach toward target locations in constrained, lateral access scenarios by employing motion primitives to react to contacts[31]. All of these approaches assume that target locations are known a priori and do not support target object extraction. Blind object recognition and retraction have been demonstrated using both non-prehensile tactile sensing[32] and prehensile tactile sensing[33], but these works do not investigate the interplay of vision and force sensing.
@@ III-A Hardware
In addition to camera and tactile information, we provide proprioception readings using the robot’s joint encoders to form the tool coordinate pose (TCP), which is located at the suction cup and consists of the cartesian position concatenated with the quaternion orientation. We also use built-in Flexiv commands that leverage joint-torque sensing to determine the dynamics-compensated wrench, represented in the TCP frame. Considering the force components of this wrench (moments are small at the suction cup), the minimum signal change that can reliably be detected above noise due to motion of the arm is $\approx 3.3\,N$.
@@ III-B Training Setup
Fig. 3 shows the training and inference pipeline used in our system. The $128\times 128$ pixel “eye-in-hand” camera images are fed through a designated ResNet-18 vision encoder, while the $20\times 5$ tactile images are fed through a separate ResNet-18 encoder. We choose ResNet since it was found to be the most effective pre-trained visual encoder for use on tactile information in [5] and its effectiveness was validated in [37]. After normalization, pose, pressure, and wrench are concatenated and fed as separate, low-dimensional inputs to the diffusion policy framework. Diffusion policy then outputs actions that can be executed by the robot controller. The action space at timestep $i$ is an 8-dimensional vector of the form
$\color[rgb]{0,0,0}A_{i}=[x_{i},y_{i},z_{i},q_{x,i},q_{y,i},q_{z,i},q_{w,i},g_{i}]$ (1)
where $x_{i},y_{i},z_{i}]$ is the commanded cartesian position, $q_{x,i},q_{y,i},q_{z,i},q_{w,i}]$ is the commanded quaternion orientation, and $g_{i}$ is the binary gripper command that controls suction.
@@ IV-B Task Setup
The task involves starting at a home position in front of the cabinet shelf with a goal of gently navigating through the cluttered environment to acquire a target object and bring it out of the shelf. Successful completion of the task requires locating the correct target object (the only red object in the scene), retracting it to the home position, and dropping it into the region outlined in black (visible beneath the suction cup in the foreground of Fig. 4c,d). We require that the task is completed within 120 seconds. Since we use a sampling rate of $f_{s}=10\,Hz$, timeout failures will result in an episode length that contains $\approx 1,200$ action steps, whereas successful trials will fall into a distribution of shorter episodes.
@@
where $\delta t_{react}$ is the time given to both demonstrator and policy to react to the contact and $F_{net}$ is the L-2 norm of the first 3 elements of the measured wrench averaged over the previous duration of $\delta t_{react}$. We also require that motions cannot exceed a peak force impulse, $I_{peak}$
@@
where $F_{peak}$ is the maximum L-2 norm of the triaxial tactile forces across all taxels. This value of $F_{peak}$ is also averaged over the previous duration of $\delta t_{react}$ at each timestep.
We implement two distinct impulse thresholds because the peak force is the metric most likely to damage an object and the wrench measurements are unaware of individual contact magnitudes when there are multiple simultaneous contacts. The peak tactile force, however, does not have full coverage of the robot arm and will be unaware of some contacts, thus a net force threshold is also needed.
@@
Through this process, we establish the excessive impulse task constraints of $I_{net}<20.8\,Ns$ and $I_{peak}<4.8\,Ns$. Adjusting these thresholds for alternative tasks merely requires repeating these initial safety tests on the new scene.
@@ IV-C Demonstration Setup
Using a 3Dconnexion SpaceMouse® Compact to teleoperate the robot, we collect demonstrations on randomized cluttered scenes generated through the process described in Section IV-A. The teleoperation is performed by mapping the $x$, $y$, $z$, and yawyaw signals of the SpaceMouse to incrememental changes in the $x$, $y$, $z$, and $\theta_{z}$ values of the robot’s commanded TCP. The $\theta_{x}$ and $\theta_{y}$ values are fixed to maintain alignment with the cabinet. The left button of the SpaceMouse turned suction on by switching a 3-way solenoid valve to the vacuum pump, whereas the right button switched the solenoid valve to atmospheric pressure, quickly releasing the object. As seen in Fig. 5, the demonstrator has access to the “eye-in-hand” camera view, the tactile images as described in Fig. 2, and pressure sensing, as well as an additional “excessive force” image which visually encodes whether the robot is approaching $F_{net}$ and $F_{peak}$. The left and right sides of this excessive force image (shown in Fig. 5b) represent peak tactile force from the left and right sides of the end-effector while the center portion corresponds to the net force magnitude computed using the joint torques. In these portions, we display grayscale values ranging from black to white as the corresponding force values range from 0% to 8080% of $F_{net}$ and $F_{peak}$. Beyond 8080%, we display a red warning in the corresponding portion to enable the demonstrator to respond. This visualization system was empirically tuned to provide the demonstrator with enough warning to avoid excessive force failures while avoiding false positives as necessary non-prehensile maneuvers are performed.
To maintain a consistent demonstrated policy that spans a range of strategies, the following criteria were used to determine actions during teleoperated task completion. If the operator saw the target object to start the task, progress directly toward the target was commanded. If the target object was not initially visible, each task would alternate whether initial progress was commanded forward and left (favoring the +y+y direction) or forward and right (favoring the y-y direction). Once the target became visible, the teleoperator would resume progress toward the target object. While making progress directly toward the target object, the operator pushed impeding obstacles aside to clear space, resulting in a roughly sinusoidal, snaking motion. During these pushing maneuvers, if obstacles jammed against the environment constraints, the excessive force indicator warned the teleoperator to move away from the jammed cluster of obstacles and reorient to maintain focus on the target object. If object jamming persisted with no clear path remaining toward the target, the operator retracted to the opening of the lateral access scene and re-entered a new location to break apart the jammed cluster of obstacles. Once the target was isolated, the demonstrator acquired the target object and retracted it to the final goal position.
@@ V-A Force Ablation Experiment
With these demonstrations and the architecture described in Section III-B, we trained four policies with varied access to wrench and tactile information to investigate their importance for this task: 1) the baseline policy masks out (replaces with zero values) wrench and tactile information, 2) the wrench-informed policy masks out tactile information, 3) the tactile-informed policy masks out wrench information, and 4) the wrench + tactile policy has both wrench and tactile information included. Apart from this masking, all policies are trained using the same architecture and hyper-parameters, with each policy trained for $200$ epochs.
@@ V-B Results
The results of $40$ evaluations on each policy are summarized in Fig. 6. The first item of note is that each policy that has access to wrench or tactile information in addition to vision significantly outperforms the baseline policy. This can be seen in Fig. 6a, which displays the number of successes, time-out failures, and excessive force failures for each policy. The best performance is demonstrated by the wrench + tactile policy, where we see an 8080% relative improvement above the baseline. The improved performance for all policies with access to either force modality appears to be largely due to a substantial reduction to excessive force failures.
@@
Figure 6: Performance results for 40 evaluations of the force ablation experiment. a) Proportions of successes, time out failures, and excessive force failures for each of the policies. * indicates $p<0.05$ in a two-sided pairwise z-test against the baseline. indicates $p<0.05$ in a two-sided pairwise z-test against the wrench + tactile policy. b) Success times for each policy with error bars showing standard error. Each policy with any force modality improved success rate and accomplished the task faster.
@@
The wrench + tactile policy appears to derive its improved performance largely from the ability to avoid time-out failures as seen in Fig. 6a. In contrast, the significantly higher proportion of timeout failures for the remaining policies often occurred due to pulling away from objects in the absence of large contact forces, failing to focus on the target after it has been isolated, and even retracting without the target object. These behaviors are shown in the accompanying video in the supplementary materials. The baseline policy exhibits the highest timeout failure rate. Finally, as seen in Fig. 6b, all policies informed by wrench or tactile information accomplished the task faster, with the wrench-informed policy being fastest.
@@
Figure 7: Heatmaps of resulting motion after measured net and peak forces for all policies. a) Motion performed by the baseline policy as a function of measured net forces. b) Resulting motion versus measured peak forces for the baseline policy. c) Resulting motion as a function of measured net force for the wrench-informed policy. d) Resulting motion as a function of measured peak force for the wrench-informed policy. e) Resulting motion as a function of measured net force for the tactile-informed policy. f) Resulting motion as a function of measured peak force for the tactile-informed policy. g) Resulting motion as a function of measured net force for the wrench + tactile policy. h) Resulting motion as a function of measured peak force for the wrench + tactile policy. The black dashed lines represent the excessive force values used for our task. The baseline policy displays the highest proportion of large force, low motion datapoints, indicating a failure to react to contacts.
@@
To gain a better understanding of how the policies have learned to react to contact forces, we extract the measured forces and resulting displacements of the end-effector for all timesteps of all evaluations for each policy. We compute the resulting displacement over a subsequent duration of 0.8s0.8\,s to allow for 1 full inference cycle to enable proper reactions.
@@
The observed motion versus net forces and peak tactile forces can be seen in Fig. 7, where cells toward the bottom right of each plot correspond to a large observed force but little resulting motion—indicating a failure to respond. For reference, the values of $F_{net}$ and $F_{peak}$ are plotted as vertical dotted lines. Ideally, there should be no cells that correspond to little or no motion when forces exceed or approach these values, since remaining in place would maintain the excessive force and cause an impulse failure. As seen in Fig. 7a,b, the baseline policy contains a substantial number of high force, low motion datapoints, but the remaining policies in Fig. 7c-h largely avoid low motion as the observed forces increase. Notably, the tactile-informed policy appears to have more datapoints at large peak force and low motion than either the wrench-informed or wrench + tactile policies, indicating that wrench sensing appears to be the most effective modality for reacting to large forces.
@@
We also note the number of failed reactions by tallying the total instances where less than 0.5cm0.5\,cm of displacement is seen after observing either a net force that exceeds $F_{net}$ or a peak tactile force that exceeds $F_{peak}$. The baseline policy fails to react to large contact forces $412$ times in contrast to $17$ failed reactions for the wrench-informed policy, $72$ for the tactile-informed policy, and $22$ for the wrench + tactile policy.
@@ VI Discussion
When attempting to accomplish object retraction in constrained clutter using imitation learning, we have demonstrated that the addition of force sensing modalities intuitively improves a policy’s ability to be gentle. Beyond this, we note that each policy in our experiment which lacked any force modality that the demonstrator used to react saw a significant increase in timeout failures, during which the policies displayed erratic behavior. This may indicate that the performance improvements seen with the addition of contact force sensing result from enabling policies to correctly identify the cause-effect relationship between observations and expert actions. Previous work establishes a framework to interpret causal relationships in imitation learning by supposing that there are disentangled factors of the observation data, some of which are true causes for expert actions and some of which are nuisance factors that coincidentally correlate with expert actions[38]. This prior work demonstrates that the removal of nuisance factors from the observation data can reduce causal confusion in imitation learning. Since highly cluttered and occluded environments have many visual distractors that do not correlate to expert actions, our experiments suggest that the addition of true causes of expert actions—in our case, contact force sensing—may be an alternative method of reducing causal confusion amidst visual distractors, though future experimentation is needed to verify this correlation.
It is worth noting that the wrench + tactile policy exhibited a higher excess force failure rate than the wrench-informed and tactile-informed policies. While this counter-intuitive result is not statistically significant and may be a result of chance, an additional explanation could be that policies containing more observation data modalities are often less data efficient, requiring more training data to achieve satisfactory performance [39, 40]. Since the number of demonstrations used in our study is relatively low, the wrench + tactile policy may suffer from excess force failures due to this data efficiency issue.